OpenStudy (anonymous):

If x^2+3<4 and 2x^2+3>4, what is a possible value of x?

OpenStudy (anonymous):

x^2+3<4 and 2x^2+3>4 \[ x^2 < 1\\ -1< x < 1\\ 2x^2+3>4\\ 2 x^2 > 1\\ x^2> 1/2\\ x< -\frac{\sqrt 2}2 ,\quad x>\frac{\sqrt 2}2 \] What's for both?

OpenStudy (anonymous):

x<-1 and x<1

OpenStudy (anonymous):

\[ -1< x < -\frac {\sqrt 2}2 \quad \text{ or } \quad \frac {\sqrt 2}2 < x < 1 \]

OpenStudy (anonymous):

so it can be -1 or 1, right?

OpenStudy (anonymous):

No if you have strict inequality to begin with.

OpenStudy (anonymous):

i dont think we do here

OpenStudy (anonymous):

So the answer is \[ -1< x < -\frac {\sqrt 2}2 \quad \text{ or } \quad \frac {\sqrt 2}2 < x < 1 \]